# Regularization in Data-driven Predictive Control: A Convex Relaxation Perspective

### v1

@@ Abstract
This paper explores the role of regularization in data-driven predictive control (DDPC) through the lens of convex relaxation, in line with [1]. Using a bi-level optimization framework, we model system identification as an inner problem and predictive control as an outer problem. Within this framework, we show that several regularized DDPC formulations, including $\ell_{1}$-norm penalties, projection-based regularizers, and a newly introduced causality-based regularizer, can be viewed as convex relaxations of their respective bi-level problems. This perspective clarifies the conceptual links between direct and indirect data-driven control and highlights how regularization implicitly enforces system identification. We further propose an optimality-based variant, O-DDPC, which approximately solves the inner problem with all identification constraints via an iterative algorithm. Numerical experiments demonstrate that O-DDPC outperforms existing regularized DDPC by reducing both bias and variance errors. These results indicate that further benefits may be obtained by applying system identification techniques to pre-process the trajectory library in nonlinear settings. Overall, our analysis contributes to a unified convex relaxation view of regularization in DDPC and sheds light on its strong empirical performance beyond linear time-invariant systems.
@@
thanks: This work is supported by NSF ECCS-2154650, CMMI-2320697, CAREER 2340713, and an Early Career Faculty Development Award from the Jacobs School of Engineering at UC San Diego. The material in this paper was partially presented at the 6th Annual Learning for Dynamics & Control Conference, University of Oxford, 15-17 July, 2024 [2].
@@ 1 Introduction
There has been a surging interest in utilizing data-driven techniques to control systems with unknown dynamics [3, 4, 5, 6]. Existing data-driven methods can be generally categorized into indirect and direct control techniques. The indirect data-driven control approaches typically include the sequential system identification (system ID) and model-based control [7, 8, 9, 10]. This two-stage control pipeline has been widely used, especially for linear systems. More recently, the Koopman operator has been leveraged to construct models of unknown nonlinear systems [11, 12, 13], but the accuracy of such models highly depends on the choice of lifting functions that are non-trivial to select [14]. In contrast, direct data-driven control methods bypass system identification altogether and design control strategies directly from input-output data, offering practitioners a potentially more attractive end-to-end alternative [4, 1, 6].
@@
One popular direct approach is the data-driven predictive control (DDPC) [4], which utilizes Willem’s fundamental lemma [15] to construct a data-driven representation of the system and use it in receding horizon control. One of the earliest formulations is the so-called DeePC, proposed in [16]. DeePC is initially established for deterministic linear time-invariant (LTI) systems, and its equivalence with subspace predictive control (SPC) has been discussed in [17]. Subsequent works [18, 19] have further investigated terminal constraint design for the closed-loop stability in LTI systems.
@@
DeePC and its general DDPC variants have demonstrated promising experimental results for controlling systems beyond LTI settings [20, 21, 22, 23, 24]. The recent work [25] has established some theoretical guarantees for DDPC in nonlinear systems. For non-deterministic or nonlinear systems, both offline and online collected data are needed to increase the width (i.e., column number) and depth (i.e., row number) of the trajectory library, so that an accurate enough data-driven representation can be constructed. The benefits of increasing its width are well-recognized in the literature [26, 21, 22, 23], and the recent works have further emphasized the importance of enlarging its depth [14, 27]. Moreover, appropriate regularizations are very important for DDPC to ensure reliable closed-loop performance [28, 1, 4].
@@
Figure 1: Schematic of data-driven predictive control (DDPC), which starts by collecting data (usually noisy) from the real system. Indirect methods identify a parametric model, while DDPC forms a Hankel matrix as the trajectory library for predictive control. Our bi-level formulation (7) integrates system ID techniques for trajectory library in DDPC. We introduce a series of convex relaxation (23), (29), (33) and approximation (38) for the bi-level formulation.
@@
Since the introduction of DeePC in [16], numerous regularization strategies for DDPC have been developed [28, 1, 29, 30, 21]. Early approaches based on $\ell_{1}$- and $\ell_{2}$-norm penalties were primarily heuristic, which aim to improve empirical control performance and ensure numerical stability for systems beyond LTI settings [16, 29, 30, 21]. A novel projection-based regularizer was later proposed in [1], establishing a formal connection between DeePC and SPC; in fact, the two formulations are equivalent when the weighting parameter is sufficiently large. Another line of work introduced $\gamma$-DDPC [28], which reformulates the original data-driven representation via an LQ factorization, introducing a new decision variable $\gamma$. Regularization is then applied directly to $\gamma$, which offers a potentially transparent interpretation of the regularization effects. Moreover, it is shown in [31, Theorem 2] that $\gamma$-DDPC with $\ell_{2}$ regularization and a sufficiently large weight is equivalent to SPC. This framework was further extended in [31] by incorporating causality constraints and designing an associated regularizer. More recently, [32] proposed a maximum-likelihood estimator that characterizes future input-output trajectories through an iterative algorithm, thereby eliminating the need for manual parameter selection.
@@
The advantages and limitations of direct and indirect data-driven methods have been extensively discussed; see the editorial column [6] for an excellent overview. Recent works have further examined the role of regularization in connecting these two paradigms. A notable contribution is [1], which analyzes the effect of regularization in a principled way via a bi-level optimization framework. In this setting, indirect data-driven control is formulated as a bi-level problem involving both identification and control, and many regularized variants of DDPC (e.g., $\ell_{1}$-norm penalties and projection-based regularizers) can be interpreted as convex relaxations of this formulation. A similar analysis has been carried out for $\gamma$-DDPC in [31], where the identification task serves as the inner problem. An alternative perspective is established in [33], which interprets the regularization term as an implicit predictor: the regularizer implicitly selects a model class for the data-driven representation and reduces model complexity. The recent work [34] introduces the concept of the final control error (FCE), defined as the expected control cost with respect to the model distribution given pre-collected data. The proposed FCE-DDPC minimizes this measure and includes certain regularized DDPC and $\gamma$-DDPC as suboptimal instances [34].
@@
In this paper, we adopt a bi-level optimization framework, similar to [1], to study the interplay between direct and indirect data-driven control. In this formulation, data pre-processing is modeled as an inner optimization problem (identification), while predictive control is formulated as an outer optimization problem (online control). Figure 1 illustrates the overall process. Our main technical contributions are as follows.
@@
First, we highlight the role of the Hankel trajectory library $H$ as a non-parametric model within the bi-level framework. This choice not only facilitates the incorporation of common system ID constraints (e.g., SPC [35], low-rank approximation [36], and causal models [37]) into the inner problem, but also enables the use of penalty methods to reduce the bi-level formulation to a single-level problem (Theorem 1). Second, under this bi-level framework, we establish three regularized DDPC formulations as convex relaxations of their corresponding bi-level optimization problems (Theorems 2, 3, 4). In each case, the explicit projection of data onto LTI identification constraints is replaced with suitable regularizations that account for implicit identification. While this perspective aligns with [1], we refine the analysis by clarifying conceptual inaccuracies, providing elementary self-contained proofs, and introducing a new causality-based regularizer (Theorem 3). Finally, all three regularized DDPC formulations only implicitly enforce a subset of system ID constraints while neglecting others. To address this limitation, we propose an optimality-based variant of DDPC, O-DDPC, which approximately solves the inner identification problem with all system ID constraints via an iterative algorithm. This leads to a refined data-driven system representation that captures linearity, causality, and the dominant trajectory space. Numerical experiments demonstrate that O-DDPC outperforms existing regularized DDPC approaches in reducing both bias and variance errors.
@@
The rest of this paper is structured as follows. Section 2 reviews the preliminaries and introduces the problem statement. Section 3 discusses the use of the trajectory matrix as a predictive model in both indirect and direct approaches. In Section 4, we present three regularized DDPC formulations as convex relaxations of suitable bi-level problems. Section 5 develops the proposed optimality-based DDPC. Numerical results are reported in Section 6, and Section 7 concludes the paper. Some technical proofs are provided in the appendix.
@@ 2 Preliminaries
This section reviews model-based predictive control, the fundamental Lemma [15] from a behavioral perspective, and a basic data-driven predictive control in [16].
@@ 2.1 LTI systems and model predictive control
$x(k+1)$ $=Ax(k)+Bu(k),$ (1)
$y(k)$ $=Cx(k)+Du(k),$
where the state, input, output at time $k$ are $x(k)\in\mathbb{R}^{n}$, $u(k)\in\mathbb{R}^{m}$, and $y(k)\in\mathbb{R}^{p}$, respectively. Throughout this paper, we assume that $(A,B)$ is controllable and $(C,A)$ is observable. The lag of the LTI system (1) is defined as the minimum integer $l\in\mathbb{N}$ such that its observability matrix $\textnormal{col}(C,CA,\ldots,CA^{l-1})$ has full column rank $n$. It is known that $l\leq n$ when $(C,A)$ is observable.
@@
$\min_{\bar{x},\bar{u},\bar{y}}\;\,$ $\sum_{k=t}^{t+N-1}\big{(}\|\bar{y}(k)-y_{\textrm{r}}(k)\|_{Q}^{2}+\|\bar{u}(k)\|_{R}^{2}\big{)}$
$\mathrm{subject~to}\;\,$ $\bar{x}(k+1)=A\bar{x}(k)+B\bar{u}(k),$ (2a)
$\bar{y}(k)={C}\bar{x}(k)+{D}\bar{u}(k),$ (2b)
$\bar{x}(t)=x_{t},$ (2c)
$\bar{u}(k)\!\in\!\mathcal{U},\bar{y}(k)\!\in\!\mathcal{Y},k=t,\ldots,t\!+\!N\!-\!1,$ (2d)
where $x_{t}\!\in\!\mathbb{R}^{n}$ is the initial state at time $t$, $\bar{x}(k),\bar{u}(k)$ and $\bar{y}(k)$ denote the predicted state, input, and output at time $k$, and $R$ and $Q$ are positive definite cost matrices. According to (2a)-(2c), $\{\bar{x}(k),\bar{u}(k),\bar{y}(k)\}_{k=t}^{t+N-1}$ is a predicted trajectory of the system (1) with an initial state $x_{t}$ at time $t$ using the system model (1). Equation (2d) enforces the constraints on the predicted trajectory, and we assume $\mathcal{U}$ and $\mathcal{Y}$ are convex sets. To simplify notation and without loss of generality, we consider a regulation problem (i.e., $y_{\textrm{r}}=\mathbb{0}_{pN}$) for the rest of the discussions.
As commonly used in MPC, the first optimal input of (2) is applied to the system (1) and a new optimal input is computed based on a new state measurement at the next step $t+1$. Guarantees on closed-loop performance (e.g., closed-loop stability and constraint satisfaction) can be ensured either by a sufficiently long horizon $L$ or by incorporating suitable terminal ingredients [38]. It is clear that (2) is a convex optimization problem (it is a quadratic program when $\mathcal{U}$ and $\mathcal{Y}$ are polytope), which admits an efficient solution when the model for (1) is known, i.e., matrices $A$, $B$, $C$ and $D$ are known.
@@
Then, problem (2) can be implemented by either indirect system identification and model-based control [39, 38] or the recent direct data-driven predictive control, such as the so-called DeePC [16] and its related approaches [1, 4, 28]. The advantages and limitations of these two classes of methods have been extensively discussed in the literature; see, for example, the recent editorial column in [6].

### v2

@@ Abstract
This paper explores the role of regularization in data-driven predictive control (DDPC) through the lens of convex relaxation, in line with [dorfler2022bridging]. Using a bi-level optimization framework, we model system identification as an inner problem and predictive control as an outer problem. Within this framework, we show that several regularized DDPC formulations, including $l_{1}$-norm penalties, projection-based regularizers, and a newly introduced causality-based regularizer, can be viewed as convex relaxations of their respective bi-level problems. This perspective clarifies the conceptual links between direct and indirect data-driven control and highlights how regularization implicitly enforces system identification. We further propose an approximation-based variant, A-DDPC, which approximately solves the inner problem with all identification constraints via an iterative algorithm. Numerical experiments demonstrate that A-DDPC outperforms existing regularized DDPC by reducing both bias and variance errors. These results indicate that further benefits may be obtained by applying system identification techniques to pre-process the trajectory library in nonlinear settings. Overall, our analysis contributes to a unified convex relaxation view of regularization in DDPC and sheds light on its strong empirical performance beyond linear time-invariant systems.
@@
thanks: This work is supported by NSF ECCS-2154650, CMMI-2320697, CAREER 2340713, and an Early Career Faculty Development Award from the Jacobs School of Engineering at UC San Diego. The material in this paper was partially presented at the 6th Annual Learning for Dynamics & Control Conference, University of Oxford, 15-17 July, 2024 [shang2024convex].
@@ 1 Introduction
There has been a surging interest in utilizing data-driven techniques to control systems with unknown dynamics [hu2023toward, markovsky2021behavioral, talebi2024policy, dorfler2023data]. Existing data-driven methods can be generally categorized into indirect and direct control techniques. The indirect data-driven control approaches typically include the sequential system identification (system ID) and model-based control [ljung1998system, chiuso2019system, zheng2021sample, kouvaritakis2016model]. This two-stage control pipeline has been widely used, especially for linear systems. More recently, the Koopman operator has been leveraged to construct models of unknown nonlinear systems [korda2018linear, haseli2023modeling, mauroy2020koopman], but the accuracy of such models highly depends on the choice of lifting functions that are non-trivial to select [shang2024willems]. In contrast, direct data-driven control methods bypass system identification altogether and design control strategies directly from input-output data, offering practitioners a potentially more attractive end-to-end alternative [markovsky2021behavioral, dorfler2022bridging, dorfler2023data].
@@
One popular direct approach is the data-driven predictive control (DDPC) [markovsky2021behavioral], which utilizes Willems’ fundamental lemma [willems2005note] to construct a data-driven representation of the system and use it in receding horizon control. Early ideas along this line can be traced back to the predictive control approach based on reduced Hankel matrices [yang2015data]. Later, one of the most influential formulations, the so-called DeePC, is proposed in [coulson2019data]. DeePC is initially established for deterministic linear time-invariant (LTI) systems, and its equivalence with subspace predictive control (SPC) has been discussed in [fiedler2021relationship]. Subsequent works [berberich2020data, berberich2021design] have further investigated terminal constraint design for the closed-loop stability in LTI systems. Efficient DDPC formulations for LTI systems have been developed in [alsalti2024robust, o2022data], which require less offline data and can handle potentially noisy and irregular data. For systems beyond LTI settings, DDPC is extended to linear parameter varying systems [verhoek2021data] and the feedback linearizable nonlinear systems [alsalti2023data]. The work [lian2021nonlinear] utilizes reproducing kernel functions to lift the state and obtain a linear model in the reproducing kernel Hilbert space, while the work [lazar2024basis] uses basis functions to synthesize the data-driven representation in DDPC.
@@
DeePC and its general DDPC variants have demonstrated promising experimental results for controlling systems beyond LTI settings in various real-world applications, e.g., robotics [elokda2021data, fawcett2022distributed], mixed traffic systems [wang2023implementation, Shang2025Decentralized], power systems [lian2023adaptive, huang2021decentralized]. The recent work [berberich2022linear] has established some theoretical guarantees for DDPC in nonlinear systems. For non-deterministic or nonlinear systems, more offline and online collected data are needed to increase the width (i.e., column number) and depth (i.e., row number) of the trajectory library, so that an accurate enough data-driven representation can be constructed. The benefits of increasing its width are well-recognized in the literature [zhang2023dimension, wang2023deep, lian2023adaptive, wang2023implementation], and the recent works have further emphasized the importance of enlarging its depth [shang2024willems, lawrence2024deep].
@@
Figure 1: Schematic of data-driven predictive control (DDPC), which starts by collecting data (usually noisy) from the real system. Indirect methods identify a parametric model, while DDPC forms a Hankel matrix as the trajectory library for predictive control. The bi-level formulation (7) integrates system ID techniques for trajectory library in DDPC. We introduce a series of convex relaxation (23), (29), (33) and an approximation (38) for the bi-level formulation.
@@
Since the original DeePC in [coulson2019data], numerous regularization strategies for DDPC have been developed [breschi2023data, dorfler2022bridging, huang2021quadratic, huang2021decentralized, wang2023deep, sader2023causality, yin2021maximum]. Early approaches based on $l_{1}$- and $l_{2}$-norm penalties were primarily heuristic, which aim to improve empirical control performance and ensure numerical stability for systems beyond LTI settings [coulson2019data, huang2021quadratic, huang2021decentralized, wang2023deep]. A novel projection-based regularizer was later proposed in [dorfler2022bridging], establishing a formal connection between DeePC and SPC; in fact, the two formulations are equivalent when the weighting parameter is sufficiently large. Another line of work introduced $\gamma$-DDPC [breschi2023data], which reformulates the original data-driven representation via an LQ factorization and introduces a new decision variable $\gamma$. Regularization is then applied directly to $\gamma$, which offers a potentially transparent interpretation of the regularization effects. Moreover, it is shown in [sader2023causality, Theorem 2] that $\gamma$-DDPC with $l_{2}$ regularization and a sufficiently large weight is equivalent to SPC. This framework was further extended in [sader2023causality] by incorporating causality constraints and designing an associated regularizer. More recently, [yin2021maximum] proposed a maximum-likelihood estimator that characterizes future input-output trajectories through an iterative algorithm.
@@
The advantages and limitations of direct and indirect data-driven methods have been extensively discussed; see the editorial column [dorfler2023data] for an excellent overview. Recent works have further examined the role of regularization in connecting these two paradigms. A notable contribution is [dorfler2022bridging], which analyzes the effect of regularization in a principled way via a bi-level optimization framework. In this setting, indirect data-driven control is formulated as a bi-level problem involving both identification and control, and many regularized variants of DDPC (e.g., $l_{1}$-norm penalties and projection-based regularizers) can be interpreted as convex relaxations of this formulation. A similar analysis has been carried out for $\gamma$-DDPC in [sader2023causality], where the identification task serves as the inner problem. An alternative perspective is established in [kladtke2023implicit], which interprets the regularization term as an implicit predictor: the regularizer implicitly selects a model class for the data-driven representation. The recent work [chiuso2023harnessing] introduces the concept of the final control error (FCE). The proposed FCE-DDPC minimizes this measure and includes certain regularized DDPC and $\gamma$-DDPC as suboptimal instances [chiuso2023harnessing].
@@
In this paper, we adopt a bi-level optimization framework, similar to [dorfler2022bridging]. We aim to study the interplay between direct and indirect data-driven control for systems beyond LTI settings, especially for the effect of regularizers. In this formulation, data preprocessing is modeled as an inner optimization problem (identification), while predictive control is formulated as an outer optimization problem (online control). Figure 1 illustrates the overall process. Our main technical contributions are as follows.
@@
First, we highlight the role of the Hankel trajectory library $H$ as a non-parametric model within the bi-level framework. This choice not only facilitates the incorporation of common system ID constraints (e.g., SPC [favoreel1999spc], low-rank approximation [markovsky2016missing], and causal models [qin2005novel]) into the inner problem, but also enables the use of penalty methods to reduce the bi-level formulation to a single-level problem (Theorem 1). Second, under this bi-level framework, we establish three regularized DDPC formulations as convex relaxations of their corresponding bi-level optimization problems (Theorems 2, 3, 4). In each case, the explicit projection of data onto LTI identification constraints is replaced with suitable regularizations that account for implicit identification. While this perspective aligns with [dorfler2022bridging], we refine the analysis by clarifying conceptual inaccuracies, providing elementary self-contained proofs, and introducing a new causality-based regularizer (Theorem 3). Finally, all three regularized DDPC formulations only implicitly enforce a subset of system ID constraints while neglecting others. To address this limitation, we propose an approximation-based variant of DDPC (A-DDPC), which approximately solves the inner identification problem with all system ID constraints via an iterative algorithm. This leads to a refined data-driven representation that approximately captures linearity, causality, and the dominant trajectory space. Numerical experiments show that A-DDPC outperforms existing regularized DDPC approaches in reducing both bias and variance errors.
Some preliminary results appeared in [shang2024convex]. In this work, we make several substantial new contributions: 1) we prove that the regularized DDPC formulations are convex relaxations of specific bi-level optimization problems and the associated regularizers are exact penalties, rather than merely convex approximations as discussed in [shang2024convex]; 2) we explicitly introduce the causality requirement and systematically integrate it into the proposed bi-level optimization framework; and 3) we enhance numerical experiments for systems beyond the LTI setting. Finally, we have included the theoretical proofs omitted in [shang2024convex].
@@
The rest of this paper is structured as follows. Section 2 reviews the preliminaries. Section 3 discusses the use of the trajectory matrix as a predictive model in indirect and direct approaches. In Section 4, we present three regularized DDPC formulations as convex relaxations of suitable bi-level problems. Section 5 develops an approximation-based DDPC. Numerical results are reported in Section 6, and Section 7 concludes the paper. Some technical proofs are provided in the appendix.
Notation: For a series of vectors $x_{1},\ldots,x_{n}$ and matrices $A_{1},\ldots,A_{n}$ with the same column size, we denote $\textnormal{col}(x_{1},\ldots,x_{n}):=[x_{1}^{{\mathsf{T}}},\ldots,x_{n}^{{\mathsf{T}}}]^{{\mathsf{T}}}$ and $\textnormal{col}(A_{1},\ldots,A_{n}):=[A_{1}^{{\mathsf{T}}},\ldots,A_{n}^{{\mathsf{T}}}]^{{\mathsf{T}}}$. We represent the $p$-norm of a vector $x$ as $\|x\|_{p}$ and the Frobenius norm of a matrix $A$ as $\|A\|_{F}$, and we denote $\|x\|_{X}^{2}:=x^{{\mathsf{T}}}Xx$ with a square matrix $X$. The null space and range space of $A$ are denoted by $\mathrm{Null}(A)$ and $\mathrm{Im}(A)$, respectively. We use $\mathbb{0}$ and $I$ to denote the zero matrix and identity matrix with compatible dimensions. We represent the open $p$-norm ball centered at $a\in\mathbb{R}^{n}$ with radius $r$ as $\mathcal{B}_{p}(a;r):=\{x\in\mathbb{R}^{n}\ |\ \|x-a\|_{p}<r\}$. We use $\mathcal{A}^{\circ}$ to denote the interior of the set $\mathcal{A}$.
@@ 2 Preliminaries
This section reviews model-based predictive control, the fundamental Lemma [willems2005note] from a behavioral perspective, and a basic data-driven predictive control in [coulson2019data].
@@ 2.1 LTI systems and model predictive control
$\bar{x}(k+1)$ $=A\bar{x}(k)+B\bar{u}(k),$ (1)
$\bar{y}(k)$ $=C\bar{x}(k)+D\bar{u}(k),$
where the state, input, output at time $k$ are $\bar{x}(k)\in\mathbb{R}^{n}$, $\bar{u}(k)\in\mathbb{R}^{m}$, and $\bar{y}(k)\in\mathbb{R}^{p}$, respectively. Throughout this paper, we assume that $(A,B)$ is controllable and $(A,C)$ is observable. The lag of the LTI system (1) is defined as the minimum integer $l\in\mathbb{N}$ such that its observability matrix $\textnormal{col}(C,CA,\ldots,CA^{l-1})$ has full column rank $n$. It is known that $l\leq n$ when $(A,C)$ is observable.
@@
$\min_{x,u,y}\;\,$ $\sum_{k=t}^{t+N-1}\big(\|y(k)-y_{\textrm{r}}(k)\|_{Q}^{2}+\|u(k)\|_{R}^{2}\big)$
$\mathrm{subject~to}\;\,$ $x(k+1)=Ax(k)+Bu(k),$ (2a)
$y(k)={C}x(k)+{D}u(k),$ (2b)
$x(t)=x_{t},$ (2c)
$u(k)\!\in\!\mathcal{U},y(k)\!\in\!\mathcal{Y},k=t,\ldots,t\!+\!N\!-\!1,$ (2d)
where $x_{t}\!\in\!\mathbb{R}^{n}$ is the initial state at time $t$, $x(k),u(k)$ and $y(k)$ denote the predicted state, input, and output at time $k$, and $R$ and $Q$ are positive definite cost matrices. According to (2a)-(2c), $\{x(k),u(k),y(k)\}_{k=t}^{t+N-1}$ is a predicted trajectory of the system (1) with an initial state $x_{t}$ at time $t$ using the system model (1). Equation (2d) enforces the constraints on the predicted trajectory, and we assume $\mathcal{U}$ and $\mathcal{Y}$ are convex sets. To simplify notation and without loss of generality, we consider a regulation problem (i.e., $y_{\textrm{r}}=\mathbb{0}_{pN}$) for the rest of the discussions. It is clear that (2) is a convex optimization problem (it is a quadratic program when $\mathcal{U}$ and $\mathcal{Y}$ are polytopes), which admits an efficient solution when the model for (1) is known, i.e., matrices $A$, $B$, $C$ and $D$ are known.
@@
Then, problem (2) can be implemented by either indirect system identification and model-based control [kouvaritakis2016model] or direct data-driven predictive control, such as the so-called DeePC [coulson2019data] and its related approaches [dorfler2022bridging, markovsky2021behavioral, breschi2023data]. The advantages and limitations of these two classes of methods have been extensively discussed in the literature [dorfler2023data].